OpenStudy (anonymous):

Assistance needed. Show that the following equation represents a circle x^2+y^2-4x+2y+1=0. Find the center and radius of the circle.

OpenStudy (souvik):

\[x ^{2}+y ^{2}-4x+2y+1=0\] \[(x ^{2}-4x+4)+(y ^{2}+2y+1)+1=4+1\] \[( x-2)^{2}+(y+1)^{2}=4\]

OpenStudy (souvik):

so its center is (2,-1) and radius is 2 unit

OpenStudy (anonymous):

How did you figure that?

OpenStudy (souvik):

i add 4 and 1 in both side...

OpenStudy (anonymous):

Oh, I see that now. Thank you.

OpenStudy (souvik):

you are welcome...:)